Shift report 20011009 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Oct 09 2001 - 00:16:16 EDT

  • Next message: Apache: "Shift report 20011009 00:00-08:00"

    ------------------------------------------------------------
    Date: Tue Oct  9 00:16:16 2001, Shift: 16:00-24:00
    Supervisor: J.I.Jordre, with: 
    Log book#9 pages 121 - 122
    
    Shift Summary:
    ------------------------------------------------------------
    1. About 3.5 hours of physics running (18:10 - 21:48), starting off 
       with ZDC rates about 207. We had less than half the ZDC rate of
       the other experiments. Upon request for steering MCR managed to 
       increase it 10-20%.
    2. Run 5249 with 337k trigger 1 events
       Run 5250 with 117k trigger 1 events
       I.e. finished 2M trigger 1 events with the following setting:
         MRS @ 4A & 45 degrees
         FFS @ 1/5 A field and 12 degrees
         BFS @ 0 field and 3 degrees
       Moved BFS also to 12 degrees, ready for FS @ 1/5 A field. 
       D5 switched to 4B.
    3. No new problems
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Oct 09 2001 - 00:16:54 EDT